OpenStudy (anonymous):

Find the solution to the equation 64^(4 – x) = 4^(2x)?

Parth (parthkohli):

Notice:\[64 ^{4 - x} = (4^3)^{4 - x}\]

OpenStudy (anonymous):

Oh, 4^3 is equal to 64. So would I make it: \[64^{4-x}=64^{4-x}\]

OpenStudy (lacypennelll):

:\ So look in your text book yet?

OpenStudy (anonymous):

Sadly I don't have a textbook, I do online school. I've been trying to reach my teacher all morning but she hasn't responded :/

OpenStudy (lacypennelll):

I do online school to I know how it feels when a teacher doesn't reach you back =-=

OpenStudy (anonymous):

do what @ParthKohli said rewrite \(64\) as \(4^3\) this means \(64^{4-x}=4^{3(4-x)}=4^{12-3x}\)

OpenStudy (anonymous):

then you know \(12-3x=2x\) so you can solve for \(x\)

OpenStudy (precal):

do what satellite73 tells you to do if you can create the same bases then you can set the exponents equal to each other and solve for x

OpenStudy (anonymous):

Thank you everyone for the help, I understand the problem now.

OpenStudy (precal):

yw